OpenStudy (anonymous):

The current exponential function P(x)=100*2^X measures the fish population’s growth monthly. Using complete sentences explain how to find the rate of growth every week.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

there are 4 weeks in a month .....

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

since 1x = 1 month x/4 = 1 week

OpenStudy (amistre64):

P(x)=100*2^X is per month P(x)=100*2^(X/4) is per week

OpenStudy (anonymous):

2x no?

OpenStudy (anonymous):

oh i see so so how do i explain im going to find the weekly rate

OpenStudy (amistre64):

i would try to do something like: x = 1 month, 1 week = x/4 let w = x/4 P(w) = 100*2^w = 100*2^(x/4)

OpenStudy (amistre64):

of course ive never been good at "explaining" in words :/

OpenStudy (anonymous):

Thats not a problem. :) this is another part to the question The National Fish and Wildlife Foundation has asked you to calculate when the population of the fish reaches 500.

OpenStudy (amistre64):

plug it in 500 = 100 * 2^x divide and log if need be

OpenStudy (anonymous):

ok so minus 100 right?

OpenStudy (amistre64):

divide ..

OpenStudy (amistre64):

500 = 100 * 2^x /100 /100 ---------------- 5 = 2^x

OpenStudy (anonymous):

if we divide it would be 5=2^x

OpenStudy (anonymous):

god you were faste

OpenStudy (amistre64):

lol .. now log off the 2

OpenStudy (anonymous):

log(5)=(log)2^2

OpenStudy (amistre64):

ln(5) = ln(2^x) ln(5) = x ln(2) ln(5)/ln(2) = x

OpenStudy (anonymous):

in why not log

OpenStudy (amistre64):

log is fine, works out for any base you want to apply; ln was simpler to type than log

OpenStudy (anonymous):

so that would be the final answer or i have to press equal on the calculator lol

OpenStudy (amistre64):

pressing equals would give us: x = 2.3219 months you can then rework that into days, weeks, or whatever is needed

OpenStudy (anonymous):

oh ok i can do that last part The computer system you need to input the function into only works in logarithms of base 10. Using complete sentences, explain how to convert your exponential function P(x) in a logarithmic one and then into a base 10 logarithm.

OpenStudy (amistre64):

i got no good bead on that one ... jumping in would be about as best i could accomplish log both sides and see if you can work out a P maybe?

OpenStudy (anonymous):

both sides? lol i didnt even know there were sides

OpenStudy (amistre64):

P = 100 * 2^x log(P) = log(100 * 2^x) log(P) = log(100) + log(2^x) log(P) = log(100) + x log(2) log(P) = log(10^2) + x log(2) log(P) = 2 log(10) + x log(2) log(P) = 2 + x log(2) im not sure what it means to convert it into a log

OpenStudy (amistre64):

if we wanted to find the inverse .. solving for x, we would do that process, but im not sure what it means to convert P into a log

OpenStudy (anonymous):

ok ill turn this in to the teacher and ill tell him that too! Thank You!

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

The information included in this question is the intellectual property of Florida Virtual School. Posting questions from any FLVS course and receiving answers is considered cheating.